# Self-supervised Learning of Echocardiographic Video Representations via Online Cluster Distillation

### v1

@@ Abstract
Self-supervised learning (SSL) has achieved major advances in natural images and video understanding, but challenges remain in domains like echocardiography (heart ultrasound) due to subtle anatomical structures, complex temporal dynamics, and the current lack of domain-specific pre-trained models. Existing SSL approaches such as contrastive, masked modeling, and clustering-based methods struggle with high intersample similarity, sensitivity to low PSNR inputs common in ultrasound, or aggressive augmentations that distort clinically relevant features. We present DISCOVR (Distilled Image Supervision for Cross Modal Video Representation), a self-supervised dual branch framework for cardiac ultrasound video representation learning. DISCOVR combines a clustering-based video encoder that models temporal dynamics with an online image encoder that extracts fine-grained spatial semantics. These branches are connected through a semantic cluster distillation loss that transfers anatomical knowledge from the evolving image encoder to the video encoder, enabling temporally coherent representations enriched with fine-grained semantic understanding. Evaluated on six echocardiography datasets spanning fetal, pediatric, and adult populations, DISCOVR outperforms both specialized video anomaly detection methods and state-of-the-art video-SSL baselines in zero-shot and linear probing setups, and achieves superior segmentation transfer. Code available at: https://github.com/mdivyanshu97/DISCOVR
@@ 1 Introduction
Modeling dynamic content in video data presents significant challenges due to complex spatio-temporal relationships, high redundancy between frames, and the need to capture both short- and long-range temporal dependencies wu2022self ; saha2025self . Echocardiography (heart or cardiac ultrasound) exemplifies these video understanding challenges. With high frame rates (30–80 fps), complex anatomical motion, and variability in image appearance caused by speckle, shadowing artifacts, and ultrasound probe variability (lang2015recommendations, ), automated echocardiography analysis requires sophisticated temporal modeling approaches. The information density in these videos is high, where features critical for diagnosis may appear as subtle variations in wall motion, valve function, or blood flow patterns that manifest only when viewed dynamically across multiple frames. Moreover, the appearance of the heart can change drastically across different cardiac views, patient populations, and imaging equipment. Developing robust video SSL models for comprehensive video understanding in echocardiography faces additional obstacles due to data limitations. Expert annotations are costly, labor-intensive, and if based on real-world hospital data often incomplete, capturing only specific aspects of the rich information contained in these videos. This scarcity of labeled data motivates SSL approaches that can leverage abundant unlabeled echocardiograms for model development (saha2025self, ; mishra2023dual, ; chen2024self, ).
Several SSL frameworks have been proposed for learning meaningful video representations, each with particular limitations in the echocardiography context. Masked video modeling methods (tong2022videomae, ; feichtenhofer2022masked, ; fan2023motion, ) tend to focus on reconstructing low-level image features like textures or edges, limiting their ability to capture high-level semantic information critical for clinical interpretation. This is especially problematic for ultrasound, which inherently exhibits a low signal-to-noise ratio (SNR), making approaches that rely on low-level pixel representations ineffective. Contrastive learning methods (he2020momentum, ; oord2018representation, ) struggle due to high inter-sample similarity and limited effective augmentations, making it difficult to construct informative positive and negative pairs, often leading to representation collapse. Clustering-based SSL methods have demonstrated strong semantic learning through self-distillation but rely heavily on aggressive augmentations that risk disrupting essential anatomical details required for fine-grained understanding.
@@
  • DISCOVR is, to our knowledge, the most comprehensive self-supervised video representation model for echocardiography to date. Trained solely on normal videos, it models healthy heart dynamics and detects pathology as deviations, eliminating the need for labeled abnormal cases. Evaluated across six datasets spanning fetal, pediatric, and adult cohorts, DISCOVR demonstrates strong generalization in zero shot classification, linear probing, anomaly detection, and segmentation, making it a versatile backbone for ultrasound analysis.
@@ 3.2.2 Semantic Cluster Distillation (SCD)
EchoNet- Dynamic MNAD 52.25 52.08 53.15
@@
Echo Pediatric- LVH MNAD 47.86 47.85 47.31
@@ 5 Ablation Study
Effect of Loss Components. We evaluate the effect of two core loss components used in DISCOVR: (i) the video self-distillation component ($\mathcal{L}_{\text{ssl}}^{\text{vid}}$), and (ii) the semantic cluster distillation component with online image guidance ($\mathcal{L}_{SCD}$). Table  4(a) reports the performance of these losses individually and in combination in zero-shot settings. Using only $\mathcal{L}_{\text{ssl}}^{\text{vid}}$ yields modest performance (F1 = 48.23%), as it primarily captures global temporal structure via CLS tokens but lacks guidance for fine-grained semantics. Introducing $\mathcal{L}_{SCD}$ leads to a substantial improvement (F1 = 61.45%, Balanced Accuracy = 63.20%), as the evolving image-based semantic clusters enrich the temporal features learned by the video model and encourage focus on more fine-grained, spatially grounded information.

### v2

@@ Abstract
Self-supervised learning (SSL) has achieved major advances in natural images and video understanding, but challenges remain in domains like echocardiography (heart ultrasound) due to subtle anatomical structures, complex temporal dynamics, and the current lack of domain-specific pre-trained models. Existing SSL approaches such as contrastive, masked modeling, and clustering-based methods struggle with high intersample similarity, sensitivity to low PSNR inputs common in ultrasound, or aggressive augmentations that distort clinically relevant features. We present DISCOVR (Distilled Image Supervision for Cross Modal Video Representation), a self-supervised dual branch framework for cardiac ultrasound video representation learning. DISCOVR combines a clustering-based video encoder that models temporal dynamics with an online image encoder that extracts fine-grained spatial semantics. These branches are connected through a semantic cluster distillation loss that transfers anatomical knowledge from the evolving image encoder to the video encoder, enabling temporally coherent representations enriched with fine-grained semantic understanding.Evaluated on six echocardiography datasets spanning fetal, pediatric, and adult populations, DISCOVR outperforms both specialized video anomaly detection methods and state-of-the-art video-SSL baselines in zero-shot and linear probing setups,achieving superior segmentation transfer and strong downstream performance on clinically relevant tasks such as LVEF prediction. Code available at: https://github.com/mdivyanshu97/DISCOVR
@@ 1 Introduction
Modeling dynamic content in video data presents significant challenges due to complex spatio-temporal relationships, high redundancy between frames, and the need to capture both short- and long-range temporal dependencies wu2022self; saha2025self. Echocardiography (heart or cardiac ultrasound) exemplifies these video understanding challenges mishra2024stan; mishra2025tier. With high frame rates (30–80 fps) mishra2025mcat, complex anatomical motion, and variability in image appearance caused by speckle, shadowing artifacts, and ultrasound probe variability (lang2015recommendations), automated echocardiography analysis requires sophisticated temporal modeling approaches. The information density in these videos is high, where features critical for diagnosis may appear as subtle variations in wall motion, valve function, or blood flow patterns that manifest only when viewed dynamically across multiple frames. Moreover, the appearance of the heart can change drastically across different cardiac views, patient populations, and imaging equipment. Developing robust video SSL models for comprehensive video understanding in echocardiography faces additional obstacles due to data limitations. Expert annotations are costly, labor-intensive, and if based on real-world hospital data often incomplete, capturing only specific aspects of the rich information contained in these videos. This scarcity of labeled data motivates SSL approaches that can leverage abundant unlabeled echocardiograms for model development (saha2025self; mishra2023dual; chen2024self).
Several SSL frameworks have been proposed for learning meaningful video representations, each with particular limitations in the echocardiography context. Masked video modeling methods (tong2022videomae; feichtenhofer2022masked; fan2023motion) tend to focus on reconstructing low-level image features like textures or edges, limiting their ability to capture high-level semantic information critical for clinical interpretation. This is especially problematic for ultrasound, which inherently exhibits a low signal-to-noise ratio (SNR), making approaches that rely on low-level pixel representations ineffective. Contrastive learning methods (he2020momentum; oord2018representation) struggle due to high inter-sample similarity and limited effective augmentations, making it difficult to construct informative positive and negative pairs, often leading to representation collapse. Clustering-based SSL methods have demonstrated strong semantic learning through self-distillation but rely heavily on aggressive augmentations that risk disrupting essential anatomical details required for fine-grained understanding.
@@
  • DISCOVR is, to our knowledge, the most comprehensive self-supervised video representation model for echocardiography to date. Trained solely on normal videos, it models healthy heart dynamics and detects pathology as deviations, eliminating the need for labeled abnormal cases. Evaluated across six datasets spanning fetal, pediatric, and adult cohorts, DISCOVR demonstrates strong generalization in zero-shot classification, linear probing, anomaly detection, and segmentation, and achieves state-of-the-art performance on downstream cardiac function estimation (LVEF prediction), making it a versatile backbone for ultrasound analysis.
@@ 3.2.2 Semantic Cluster Distillation (SCD)
EchoNet-
Dynamic
MNAD 52.25 52.08 53.15
@@
Echo
Pediatric-
LVH
MNAD 47.86 47.85 47.31
@@ 4.2 Segmentation Evaluation
4.3 LVEF Prediction
We evaluate the effectiveness of DISCOVR representations for downstream cardiac function estimation using the EchoNet-Dynamic ejection fraction dataset ouyang2020video. As shown in Table 4, DISCOVR achieves the lowest Mean Absolute Error (MAE) of 7.79 under the standard linear probing setup, outperforming other self-supervised baselines such as VideoMAE (8.02) and MGMAE (8.88). When fine-tuning only the last three encoder blocks, DISCOVR further reduces the MAE to 6.32, demonstrating the strength of its learned representations even with limited adaptation. In comparison, the fully supervised EchoNet-Dynamic model ouyang2020video is trained end to end with all parameters updated. Under an ejection fraction–only setup without segmentation labels, DISCOVR surpasses these fully supervised baselines, including MC3 with an MAE of 6.59, the base EchoNet-Dynamic model with 7.35, and R3D with 7.63. The full EchoNet-Dynamic architecture achieves an MAE of 4.05 using a large multi-task design with 71.1 million parameters co-trained on 20,060 manual segmentation tracings. These results show that DISCOVR, through self-supervised pretraining and partial fine-tuning, learns powerful cardiac representations that rival or exceed fully supervised models trained end to end.
Table 4: LVEF prediction results on the EchoNet-Dynamic dataset. Our self-supervised method is compared against other SSL methods and fully-supervised baselines from ouyang2020video.
Model MAE $\downarrow$ RMSE $\downarrow$ EF Labels Seg. Labels
Self-Supervised (Linear Probing)
VideoMAE 8.02 11.16
MGMAE 8.88 12.47
DISCOVR (Ours) 7.79 10.89
Self-Supervised (Fine-tuning)
DISCOVR (finetune last 3 blocks) 6.32 8.62
Fully-Supervised Baselines [1] trained only with EF Data
MC3 (All frames) 6.59 9.39
EchoNet-Dynamic (EF, All frames) 7.35 9.53
R3D (All frames) 7.63 9.75
DISCOVR (finetune last 3 blocks,64 frames) 6.32 8.62
EchoNet-Dynamic (Full model) 4.05 5.30
@@ 5 Ablation Study
Effect of Loss Components. We evaluate the effect of two core loss components used in DISCOVR: (i) the video self-distillation component ($\mathcal{L}_{\text{ssl}}^{\text{vid}}$), and (ii) the semantic cluster distillation component with online image guidance ($\mathcal{L}_{SCD}$). Table  5(a) reports the performance of these losses individually and in combination in zero-shot settings. Using only $\mathcal{L}_{\text{ssl}}^{\text{vid}}$ yields modest performance (F1 = 48.23%), as it primarily captures global temporal structure via CLS tokens but lacks guidance for fine-grained semantics. Introducing $\mathcal{L}_{SCD}$ leads to a substantial improvement (F1 = 61.45%, Balanced Accuracy = 63.20%), as the evolving image-based semantic clusters enrich the temporal features learned by the video model and encourage focus on more fine-grained, spatially grounded information. For more detailed ablation, refer to supplementary section B.1.4.
@@
Computational Cost and Scalability. We report the computational cost, of our method in Table 6. The table shows both training and inference statistics, showing GPU memory usage and F1-score for each method on EchoNet-Dynamic, for a batch size of 1, 64 frames, and a spatial size of $112\times 112$. During training, DISCOVR uses slightly more GPU memory (10.5GB) compared to prior methods (between 9.0 and 9.5GB) but achieves a notable +6.38% improvement in F1-score over the closest competitor. At inference, all methods, including DISCOVR, use identical ViViT-like encoders, resulting in nearly the same GPU memory footprint and FLOPS. This demonstrates that our method’s performance improvements come with minimal extra training cost and no penalty for inference efficiency.
Table 6: Training and inference GPU memory, FLOPS, and F1-score on EchoNet-Dynamic, batch size = 1, 16 frames, $112\times 112$ resolution.
Model Train Mem (GB) F1-score Infer Mem (GB) Infer. FLOPS
MGMAE huang2023mgmae 9.0 46.13 1.153 101.85
VideoMAE tong2022videomae 9.0 55.07 1.153 101.85
SIGMA salehi2024sigma 9.2 49.04 1.153 101.85
Video-distillation 9.5 48.23 1.153 101.85
DISCOVR (Ours) 10.5 61.45 1.153 101.85
@@ 6 Conclusion
Acknowledgments
We acknowledge financial support from InnoHK-funded Hong Kong Centre for Cerebro-cardiovascular Health Engineering (COCHE), UKRI grant EP/X040186/1, UK EPSRC grant EP/T028572/1 (VisualAI), UK EPSRC Doctoral Training Partnership award and, UKRI AIRR Early Access Project No. ANON-BYYG-VX4C-Z.
Appendix
Appendix A Dataset Distribution
This section presents the dataset distributions for our five echocardiography video datasets: FetalEcho1 (Fig.6), FetalEcho2 (Fig.7), EchoNet-Dynamic (Fig.8), EchoNet-Pediatric (Fig.9), and RVENET (Fig.10). For each dataset, the bar chart displays the number of unique samples in the training, validation, and test sets. The accompanying pie charts illustrate the class distributions (Normal vs. Abnormal) within the validation and test sets.
Figure 6: Dataset Distribution for Fetal-Echo1 dataset
Figure 7: Dataset Distribution for Fetal-Echo2 dataset
Figure 8: Dataset Distribution for Echo-Dynamic dataset
Figure 9: Dataset Distribution for Echo-Pediatric dataset
Figure 10: Dataset distribution for RVENET dataset
Appendix B Additional Results:
B.1 Full Finetuning
B.1.1 Evaluation Setup
We follow the same evaluation procedure as described in the experiments section in the main paper, but fine-tune the entire backbone along with the linear layer. All other evaluation settings remain unchanged. Results are reported on the Echonet Dynamic dataset to assess end-to-end supervised performance.
B.1.2 Evaluation Result:
Under full fine-tuning, as shown in Table 7, all models experience a drop in performance compared to their linear probing results, reflecting overfitting due to the limited labeled validation data. Despite this, DISCOVR achieves the highest F1 score of 70.44%, outperforming MGMAE (65.99%), SIGMA (61.46%), and VideoMAE (57.31%). DISCOVR’s structured representation learning, through temporal distillation and cross-modal clustering, appears to provide more robust and adaptable features, enabling it to generalize better even when fully fine-tuned on a small dataset.
Table 7: Table showing the full-finetuning result of DISCOVR compared to other baselines on the Echo-Dynamic Dataset
Model (%) Accuracy Balanced Acc. Precision Recall F1-Score
VideoMAE 57.62 57.94 58.27 57.94 57.31
SIGMA 61.69 62.00 62.41 62.00 61.46
MGMAE 65.99 66.08 66.10 66.08 65.99
Ours 70.51 70.42 70.50 70.42 70.44
B.1.3 Generalisation to Other Modalities.
To test whether DISCOVR generalizes beyond echocardiography, we evaluated its transfer performance on two distinct medical image benchmarks: the Breast Ultrasound Images dataset [al2020dataset] (cancer detection across 600 patients) and DermMNIST [yang2023medmnist] (skin lesion classification). Both breast ultrasound and echocardiography require the detection of small, irregular regions of altered tissue, such as hypoechoic tumors in the breast or localized wall motion abnormalities in the heart, making the ability to identify subtle structural changes in one domain directly applicable to the other. Similarly, DermMNIST demands fine-grained visual discrimination between morphologically similar skin lesions. For both benchmarks, we froze the DISCOVR encoder and trained a linear classifier, comparing performance directly across methods.
As shown in Table 8, our method demonstrates strong generalization across both tasks. On the Breast Ultrasound dataset, DISCOVR improves balanced accuracy by 2.01% over VideoMAE, 19.83% over SIGMA, and 12.01% over MGMAE. For DermMNIST, DISCOVR achieves an accuracy of 71.68%, outperforming VideoMAE by 2.85%, SIGMA by 3.00%, and MGMAE by 3.85%. These results demonstrate strong generalization to multiple medical image analysis tasks beyond echocardiography. Further, to assess generalization to natural video data, we pretrained and evaluated all models on the Kinetics 400 action recognition benchmark, using a zero-shot protocol where KNN classification with $K=20$ was applied to features using 64 frames from the frozen video backbone. As shown in Table 9, DISCOVR achieves the highest Top-1 accuracy at 22.3%, outperforming MVD by 3.6%, MME by 3.2%, and VideoMAE by 1.6%, while also requiring the fewest pretraining epochs. These results highlight that DISCOVR not only excels at medical video tasks but also learns generalizable representations efficiently for large-scale natural video datasets.
Table 8: Linear Probing results on the Breast Ultrasound dataset and DermMNIST. For Breast Ultrasound, we report Balanced Accuracy (Bal. Acc.) and F1; for DermMNIST, we report overall Accuracy (Acc.).
Method Breast Ultrasound     DermMNIST
Balanced Accuracy F1 Score     Accuracy
VideoMAE 61.75 64.45     68.83
SIGMA 43.93 42.21     68.68
MGMAE 51.75 52.34     67.83
DISCOVR (Ours) 63.76 65.44     71.68
Table 9: Zero-Shot KNN classification performance on Kinetics-400
Model Epochs Top-1 Accuracy (%)
MVD [wang2023masked] 1600 18.7
MME [sun2023masked] 800 19.1
VideoMAE [tong2022videomae] 800 20.7
DISCOVR (Ours) 400 22.30
B.1.4 Loss function Ablation detailed.
To rigorously evaluate each component, we have added baselines using only masked image or only video self-distillation. Indeed, we find that the settings perform suboptimally, as shown in Table 1, confirming that spatial or temporal cues alone are insufficient for strong representation learning. In contrast, combining both with the SCD loss, which explicitly distills fine-grained semantic structure from the image branch into the video backbone, achieves the best results. This supports our intuition that SCD is crucial for aligning spatial semantics with temporal dynamics, enabling more robust and clinically meaningful video representations.
Table 10: Effect of the different loss terms on classification performance (Balanced Accuracy, Precision, and F1).
$\mathcal{L}^{vid}_{\text{ssl}}$ $\mathcal{L}^{img}_{\text{ssl}}$ $\mathcal{L}_{\text{SCD}}$ Bal. Acc. Precision F1
52.27 53.16 48.23
53.66 55.22 49.43
63.20 65.35 61.45
Appendix C Implementation Details
All models are implemented in PyTorch 2.6 and trained on RTX 8000 GPUs (48 GB) with a batch size of 8 using the AdamW optimizer. Videos are processed as 64-frame clips sampled at a stride of 3 and resized to $112\times 112$.
For both video and image self-distillation, we use a student-teacher setup where the teacher processes the full input and the student observes $N=4$ randomly masked views. The teacher network is updated via an exponential moving average (EMA) of the student with momentum $\lambda=0.996$. A fixed temperature $\tau_{s}=0.1$ is used for the student, while the teacher temperature $\tau_{t}$ is linearly warmed from 0.04 to 0.07 over the first 30 epochs. Semantic Cluster Distillation (SCD) uses $K=3000$ learnable prototypes, with similarity scores computed via temperature-scaled dot products ($\tau=0.1$) and cluster assignments generated using the Sinkhorn-Knopp algorithm (10 iterations, $\epsilon=0.05$). Models are trained for 400 epochs with a learning rate of $1.5\times 10^{-4}$, weight decay of 0.05, and 40 warmup epochs.
Appendix D Broader Impact and Limitations
In this work, we introduce DISCOVR, a novel self-supervised model for echocardiography video understanding across fetal, pediatric, and adult populations. Trained without labeled abnormal cases, DISCOVR learns rich spatiotemporal representations and enables zero-shot inference. One key application of DISCOVR is in the early screening of heart diseases, where it can assist clinicians by flagging potential anomalies in echocardiography videos. This has significant clinical relevance, as congenital heart defects affect approximately 1 in 100 newborns, with up to 50% missed during prenatal screening [van2011birth, knowles2014screening, van2020congenital], and cardiovascular diseases remain the leading global cause of death [who2021cvd]. By reducing reliance on large, labeled datasets, DISCOVR offers a scalable and accessible solution, particularly for deployment in low-resource settings.
While DISCOVR shows strong potential, its current scope is focused specifically on echocardiography, and it has not yet been evaluated on other imaging modalities. The model was trained and tested on five datasets collected from distinct clinical sites, each with its own imaging protocols, devices, and patient cohorts. As a result, the demographic and geographic diversity of the data may be limited. Further validation is needed to assess the model’s generalizability across broader clinical settings, populations, and imaging systems.
Figure 11: Figure (left) compares two fine-grained videos: a natural scene of a person baking (left) and an adult heart ultrasound (right). The frame-level cosine similarity matrix, computed using a pretrained VideoMAE model, shows that ultrasound frames are highly similar (mean=0.99), with only minor local variations. This highlights the difficulty in distinguishing individual frames in such medical videos. Figure (right) compares normal and abnormal fetal echocardiograms, which appear almost identical despite one being abnormal. This illustrates the inherent difficulty of distinguishing subtle cardiac abnormalities in fetal imaging.